OpenStudy (anonymous):

true or false: the ordered pair (-5,-1) is a solution to the system of equations

OpenStudy (anonymous):

false right?

OpenStudy (anonymous):

which system of equations

OpenStudy (anonymous):

x+y=4 x-y=6

OpenStudy (anonymous):

it's just plug and solve?

OpenStudy (theviper):

false

OpenStudy (anonymous):

the solution is just the x intercept so do u know how to find tht

OpenStudy (anonymous):

oh yea its false

OpenStudy (theviper):

\[\Huge{x+y=4}\]\[\Large{-5+(-1)=-5-1=4}\]\[\Large{-6\neq4}\]Similarly second ;)

OpenStudy (anonymous):

plug it in: -5+(-1)=4 -5-(-1)=6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome